OpenStudy (anonymous):

Can anyone answer this math question for me please? Simplify. (3c)^3

OpenStudy (anonymous):

27c^3

OpenStudy (sidsiddhartha):

\[(3c)^3=3^3*c^3=(3*3*3)*c^3=27c^3\] ok??

OpenStudy (anonymous):

thank you

OpenStudy (sidsiddhartha):

np :)